Adjusted WACC. Ashman Motors is currently an​ all-equity firm. It has two million shares​ outstanding, selling for $42 per share. The company has a beta of 1.4, with the current​ risk-free rate at 5.1% and the market premium at 8.7% the tax rate is 15​% for the company. Ashman has decided to sell $42 million of bonds and retire half its stock. The bonds will have a yield to maturity of 8.7​%. The beta of the company will rise to 1.9 with the new debt. What was​ Ashman's adjusted WACC before selling the​ bonds? What is its new WACC after selling the bonds and retiring the stock with the proceeds from the sale of the​ bonds?  

Hint​: The weight of equity before selling the bond is​ 100%.

What was​ Ashman's adjusted WACC before selling the​ bonds? ___%
Answer #1

Before the sale of bonds the weighted average cost of capital is the cost of equity.

rE = rF + [Beta * Market Premium]

= 5.1% + [1.4 * 8.7%] = 5.1% + 12.18% = 17.28%

After the recap.,

rE = 5.1% + [1.9 * 8.7%] = 5.1% + 16.53% = 21.63%

No. of shares to be repurchased = Amount of Debt / Current Share Price

= $42,000,000 / $42 = 1,000,000 shares

Market Value of Equity = Current Share Price * New Shares outstanding

= $42 * [2,000,000 - 1,000,000] = $42,000,000

Market Value of debt = $42,000,000

Total Market Value = Market Value of debt + Market Value of equity

= $42,000,000 + $42,000,000 = $84,000,000

WACC = [wD * kD * (1 - t)] + [wE * kE]

= [(42,000,000/84,000,000) * 8.7% * (1 - 0.15)] + [(42,000,000/84,000,000) * 21.63%]

= 3.6975% + 10.815% = 14.5125%